A raffle requires the purchase of a ticket that has 4 different numbers take from {1,2,3,...,20}. The order of the numbers does not matter. How many different raffle tickets are there?
Answer by Boreal(15235)

This is 20C4=20!/16!4!
=20*19*18*17/24
=5*19*3*17 (by canceling) =95*51=4845.
Calculator 20 MATH PRB 3 ENTER 4 ENTER